# Concentration of Measure for Distributions Generated via Diffusion Models

### v2

@@ Appendix C Gram Spectrum
Note that while establishing the closeness of eigenvalue distributions of the Gram matrices allows one to characterize the behavior of certain algorithms such as Least-Squares SVM or spectral clustering, this does not allow us to analyze more elaborate algorithms. For example, the LASSO objective $\min_{w}\|Xw-y\|+\lambda\|w\|_{1}$ for $w\in\mathbb{R}^{d},X\in\mathbb{R}^{n\times d}$ is not unitarily invariant. Hence, given $X^{\prime}\in\mathbb{R}^{n\times d}$, even knowing that the Gram matrices $(X^{\prime})^{T}X^{\prime}$ and $X^{T}X$ are exactly equal to each other does not let one conclude that $w^{\prime}$ identified via $\min_{w^{\prime}}\|X^{\prime}w^{\prime}-y\|+\lambda\|w^{\prime}\|_{1}$ yields to performance similar to the performance of $w$.
Appendix D Sampling Process
In Figures 4 and 5, we plotted the evolution of the norms. Now in Figure 13 we also observe that the sampling process first matches the higher eigenvalues of the Gram matrix, and then progressively matches the lower eigenvalues.
Figure 13: Eigenvalues of Gram matrix through EDM sampling process. Computed with 2048 samples of a single class.
D.1 Evolution of Norms of Pixels
We also investigate the norms of individual pixels through the EDM sampling process. Note that Figure 14 d is on a log-log scale, which cuts off negative values of the plotted standard deviation envelope at the low noise scales; indeed, at any step of the sampling process, there are pixels that increase in norm. But on average, the pixel norms are decreasing.
Figure 14: (a) shows distribution of the pixel norms for a single class, through the EDM sampling process. (b) shows the individual trajectories of the norms of 1000 randomly selected pixels at different noise scales of sampling. (c) shows the difference in norms between sampling steps, for 100 randomly selected pixels. (d) shows the mean and standard deviation envelope of this difference in norms. Note that this is on a log-log scale and negative values of $(\mu-\sigma)$ are cut off.
Appendix E Representations of High-Resolution Latent Diffusion Samples
@@
Figure 15: Spectra of Gram Matrices of ResNet representations of a 4-class mixture (Church, Tench, English Springer, French Horn). Computed for representations of 1350 images per class from EDM2 (Red) and for GMM fitted on those representations (Blue).
@@
As seen in Figures 15 and 16, the Gram matrices of the ResNet representations of diffusion images show a close to match to their GMM counterparts when viewing the eigenvalue spectrum. Compare with Figure 18, which presents the spectra of the covariance matrix of diffusion representations of a single class. Moreover, we observed clear separability of the classes in the first few eigenvectors. This motivated us to train a logistic regression model on the top eigenvectors of these Gram matrices, and as shown in Figure 17, the first 3-4 eigenvectors are all that are needed for near-perfect accuracy. We leave the question of how this scales with the number of classes for future work. We plot the mean and standard deviation envelope over 100 runs of logistic regression, each using a random split proportion of $0.8$ training samples from the 1350 representations per class in the mixture. The test set is fixed as a $0.2$ proportion of the representations of real images.
Figure 17: Logistic regression trained on ResNet representations of a 4-class mixture. Shown for EDM2 representations (Red), for GMM fitted on EDM2 representations (Blue), and for representations of real images (Green).
As a latent diffusion model, EDM2 Karras et al. [2024b] does diffusion in the latent space of a pre-trained variational autoencoder (VAE). We investigate the evolution of the norms of the latents through the deterministic sampling process. The results are presented in Figure 19.
Figure 18: Eigenvalues of covariance matrix of ResNet representations of EDM2 diffusion-generated images, for a single class, computed over 4096 samples.
Figure 19: Evolution of norms of latents through EDM2 deterministic sampling, for a single class. These are latents of dimension $4\times 64\times 64$ in the latent space of a pre-trained VAE. We refer to Karras et al. [2024b] for further details.
Figure 20: More samples from our diffusion-generated dataset, demonstrating the visual fidelity of the generated images. All 20 Imagenet64 classes used in our experiments are represented

### v3

@@ Appendix E Gram Spectrum
Note that while establishing the closeness of eigenvalue distributions of the Gram matrices allows one to characterize the behavior of certain algorithms such as Least-Squares SVM or spectral clustering, this does not allow us to analyze more elaborate algorithms. For example, the LASSO objective $\min_{\mathbf{w}}\|\mathbf{X}\mathbf{w}-\mathbf{y}\|_{2}^{2}+\lambda\|\mathbf{w}\|_{1}$ for $\mathbf{w}\in\mathbb{R}^{d},\mathbf{X}\in\mathbb{R}^{n\times d}$ is not unitarily invariant. Hence, given $\mathbf{X}^{\prime}\in\mathbb{R}^{n\times d}$, even knowing that the Gram matrices $(\mathbf{X}^{\prime})^{T}\mathbf{X}^{\prime}$ and $\mathbf{X}^{T}\mathbf{X}$ are exactly equal to each other does not let one conclude that $\mathbf{w}^{\prime}$ identified via $\min_{\mathbf{w}^{\prime}}\|\mathbf{X}^{\prime}\mathbf{w}^{\prime}-\mathbf{y}\|^{2}_{2}+\lambda\|\mathbf{w}^{\prime}\|_{1}$ yields performance similar to the performance of $w$.
Evolution of Gram Matrix Eigenvalues through Sampling: We observe that the sampling process first matches the top Gramian eigenvalues, and progressively matches the lower eigenvalues.
Figure 8: Gramian eigenvalues through EDM sampling. Computed with 2048 samples of one class
Appendix F More Observations
Figure 9: (A) The evolution of the $\ell_{2}$ norms through the stochastic sampling process (See algorithm 1 for definition of $D_{\theta}$ and $d_{i}$). (B) Top: The difference in $\ell_{2}$ norms of intermediate images between consecutive steps of the EDM sampling process. (B) Bottom: The mean norm decrease per class, using 2048 samples per class; shaded regions represent the variance of the norm decrease per class.
Figure 10: (A) All ordered eigenvalues ($\geq 10^{-12}$) of the covariance matrices for each class, shown log-log scale. (B) Full spectra of the covariance matrices of diffusion-generated images, computed over 10240 samples per class. Scaled by exponent $0.1$ for clearer presentation.
Appendix G Representations of High-Resolution Latent Diffusion Samples
@@
Figure 11: Spectra of Gram Matrices of ResNet representations of a 4-class mixture (Church, Tench, English Springer, French Horn) using 1350 images per class from EDM2 (Red), and for GMM fitted on those representations (Blue).
As seen in Figure 11, the Gram matrices of the ResNet representations of diffusion images show a close to match to their GMM counterparts when viewing the eigenvalue spectrum. Moreover, we observed clear separability of the classes in the first few eigenvectors. This motivated us to train a logistic regression model on the top eigenvectors of these Gram matrices, and as shown in Figure 12, the first 3-4 eigenvectors are all that are needed for near-perfect accuracy. We leave the question of how this scales with the number of classes for future work. We plot the mean and standard deviation envelope over 100 runs of logistic regression, each using a random split proportion of $0.8$ training samples from the 1350 representations per class in the mixture. The test set is fixed as a $0.2$ proportion of the representations of real images.
Figure 12: Logistic regression trained on ResNet representations of a 4-class mixture. For EDM2 representations (Red), GMM fit on EDM2 reps (Blue), reps of real images (Green).
Figure 13: Eigenvalues of single class covariance matrix of ResNet reps of EDM2 images.
@@
We also plot the top eigenspaces of the Gram matrices for multiclass mixtures of these ResNet representations, compared against GMM samples fit on those representations. Figure 14 demonstrates visually the match in the eigenvectors, and supplements the logistic regression experiments in our main body, where we trained logistic regression on the top $k$ eigenvectors of the Gram matrices and showed matching (with the GMM) test accuracies.
As a latent diffusion model, EDM2 Karras et al. [2024b] does diffusion in the latent space of a pre-trained variational autoencoder (VAE). We investigate the evolution of the norms of the latents through the deterministic sampling process, presented in Figure 15.
Figure 15: Evolution of norms of latents through EDM2 deterministic sampling, for a single class. These are latents of dimension $4\times 64\times 64$ in the latent space of a pre-trained VAE.
Appendix H Evolution of Norms of Pixels
We also investigate the norms of individual pixels through the EDM sampling process. Note that Figure 16 c is on a log-log scale, which cuts off negative values of the plotted standard deviation envelope at the low noise scales; indeed, at any step of the sampling process, there are pixels that increase in norm. But on average, the pixel norms are decreasing.
Figure 16: (a) The distribution of the pixel norms for a single class, through the EDM sampling process. (b) The individual trajectories of the norms of 1000 randomly selected pixels at different noise scales of sampling. (c) The mean and standard deviation envelope of the difference in norms of pixels between sampling steps (Note that negative values are cut off, on this log-log scale.)
Appendix I Dataset Sample
Figure 17: (A) Samples from conditional EDM diffusion model, trained on Imagenet64. (B) Distribution of $\ell_{2},\ell_{4}$ and $\ell_{10}$ norms, computed over 10240 samples per class.